A survey of 2324 adults in a certain large country aged 18 and older conducted by a reputable polling organization found that 419 have donated blood in the past two years. Complete parts (a) through (c) below. 


(a) Obtain a point estimate for the population proportion of adults in the country aged 18 and older who have donated blood in the past two years. 

(b) Verify that the requirements for constructing a confidence interval about p are satisfied. 

The sample _______ a simple random sample the value of _______ is _______ . which is _______ 10. and the _______  _______ less than or equal to 5% of the _______ 

(c) Construct and interpret a 90% confidence interval for the population proportion of adults in the country who have donated blood in the past two years. Select the correct choice below and fill in any answer boxes within your choice. (Type integers or decimals rounded to three decimal places as needed. Use ascending order.) 

A. We are _______ % confident the proportion of adults in the country aged 18 and older who have dojated blood in the past two years is between _______ and _______ . 

B. There is a _______ % probability the proportion of adults in the country aged 18 and older who have donated blood in the past two years is between _______ and _______. 
Answer #1

(a)

= 419 / 2324 = 0.180

(b)

n(1-) = 2324 * 0.180 * (1 - 0.180) = 343.0224

The sample can be assumed to be a random sample the value of n(1-) is 343.022, which is greater than or equal to 10 and the sample size can be assumed to be less than or equal to 5% of the population size.

(c)

Standard error of sample proportion, SE =

= 0.007969391

Z value for 90% confidence interval is, 1.645

90% confidence interval is,

(0.180 - 1.645 * 0.007969391 ,  0.180 + 1.645 * 0.007969391)

(0.167 ,  0.193)

A. We are 90% confident the proportion of adults in the country aged 18 and older who have donated blood in the past two years is between 0.167 and 0.193

> Thank you so much for the help! It was easy to understand.

Eaton Uter Sun, Nov 7, 2021 8:14 AM